OpenStudy (anonymous):

(1) Skip had eight fewer job interviews than Woody. (2) Together they had 20 interviews. How many interviews did each have?

OpenStudy (anonymous):

20=skip + woody skip= woody -8 20-8=12 12/2=6 6+8=14 (woody) 6(skip)

OpenStudy (anonymous):

thanks